Two hikers start walking in opposite directions from a camp at the same time. The average rate of the northbound hiker is 3 mi/h faster than the rate of the southbound hiker. If they are 22 mi apart after 2 hours, what is the rate of each hiker? (Let r be the rate of the southbound hiker)
Answer by mananth(16949)

Two hikers start walking in opposite directions from a camp at the same time. The average rate of the northbound hiker is 3 mi/h faster than the rate of the southbound hiker. If they are 22 mi apart after 2 hours, what is the rate of each hiker? (Let r be the rate of the southbound hiker)
...
rate of the southbound hiker = r mph
rate of the northbound hiker is 3 mi/h faster = x+3
..
They are movin away from each other
speed = r+r+3
=2r+3 mph.
...
distance = 22 miles
time = 2 hours
speed = 2r+3
..
t = d/r
...
2= 22/(2r+3)
2(2r+3)=22
4r+6=22
4r=22-6
4r=16
/4
r=4 mph
the other's speed = r+3 = 7mph